# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

### v3

@@ 5.1 Experimental Settings
Metrics. (1) Overall Transfer Performance: Drawing inspiration from the methodologies outlined in [17, 26], we assess the overall transfer performance by employing ArtFID [44], a metric that amalgamates LPIPS [45] for gauging content fidelity and FID [46] for evaluating the richness of style. (2) Foreground Style Preservation: To evaluate the preservation of texture, we utilize the SSIM [47] metric, while the CLIP-I score [48] is employed to assess the preservation of semantic content. During this evaluation, the background is neutralized to ensure accuracy. (3) Temporal Consistency: In line with the approaches detailed in [26, 11], we measure the similarity between consecutive frames using the CLIP-F score [48], which serves as a metric for evaluating the temporal coherence of the video. (4) Inference Efficiency: To gauge the time efficiency, we calculate the total inference time, thereby assessing their performance in terms of computational speed.
@@
5.2 Quantitative Comparison
We have chosen a selection of video stylization techniques, including Diffutoon [12] and StyleCrafter [11], as well as two video editing frameworks: AnyV2V [13] and BIVDiff [10]. Additionally, we have considered two traditional non-diffusion methods: EFDM [38] and CAST [39]. Given that our specific task does not perfectly correspond with some of the existing methods, we have made necessary modifications to ensure a fair comparison. In the case of Diffutoon [12], which is designed to process text inputs for coloring tasks and does not natively support style images, we have overcome this limitation by leveraging Chat-GPT4 [49] to generate descriptive captions for the style images. For the style generation task with StyleCrafter [11], which does not inherently support video editing, we employ inversion to obtain the initial noise of the original video, which is then utilized in the generation process. For the unified video editing frameworks, AnyV2V [13] and BIVDiff [10], we have integrated InstantStyle [50] as the image editing model to enhance their capabilities. As for the traditional non-diffusion methods, EFDM [38] and CAST [39], which are well-aligned with our task, we have retained them in their original form without making any alterations.
We employ these six established methods as benchmarks to evaluate our proposed UniVST, with the quantitative comparisons presented in Table I. The results demonstrate that our UniVST outperforms the others across all three performance metrics. On both the DAVTG-WikiArt and DAVTG-Laion datasets, UniVST consistently surpasses the competition in terms of foreground style preservation and temporal consistency, showcasing its robust capability for localized video style transfer. Regarding overall transfer performance, UniVST claims the top position on the DAVTG-Laion dataset and secures the second spot on DAVTG-WikiArt. Although localized transfer tasks might marginally affect overall performance, leading to a slightly higher FID score compared to other methods, these findings underscore UniVST’s proficiency in effectively reconciling content fidelity with style richness.
Our inference efficiency does not stand out. This is because we opt not to adopt the temporal attention mechanisms in existing T2V frameworks [11, 26]. Instead, we utilize the Sliding-Window Consistent Smoothing strategy, which requires dynamic optical flow estimation during the denoising process. While this approach improves temporal consistency, it also leads to increased computational overhead. In Fig 13, enlarging the sliding window size results in a proportional increase in the number of frames needed for optical flow estimation, thereby raising the time cost. This trade-off between temporal consistency and computational efficiency highlights the potential optimization in future work.
TABLE II: Ablation study for proposed components. The bold values represent the best performance.
Datasets Methods Overall Transfer Performance Style Preservation Consistency
@@ 5.2 Quantitative Comparison
DAVTG-WikiArt UniVST 37.152 26.134 0.369 0.986 0.991 0.980
w/o mask-guidance 38.014 26.050 0.405 0.966 0.970 0.979
w/o latent-shift 37.649 26.909 0.349 0.985 0.990 0.980
w/o attention-shift 39.589 27.506 0.388 0.985 0.988 0.983
w/o window-smoothing 38.626 26.907 0.384 0.982 0.989 0.970
DAVTG-Laion UniVST 30.636 21.146 0.383 0.986 0.990 0.981
w/o mask-guidance 30.848 20.974 0.404 0.961 0.967 0.979
w/o latent-shift 31.001 21.766 0.362 0.985 0.989 0.980
w/o attention-shift 32.534 22.623 0.377 0.985 0.989 0.981
w/o window-smoothing 30.834 21.249 0.386 0.981 0.988 0.971
@@ 5.3 Qualitative Comparison
For a qualitative assessment, we have juxtaposed the editing results with those of six other baseline methods, as illustrated in Fig. 5. Diffutoon [12], which is renowned for its expertise in cartoon coloring, exhibits a limited responsiveness to a variety of image styles. This results in a final edited video that is notably monotonous, heavily inclined towards a cartoonish anime aesthetic. StyleCrafter [11], tailored for style-based video generation, demonstrates a high sensitivity to stylistic cues. However, it encounters difficulties in preserving content fidelity, leading to substantial distortion of background details and character features during the transfer process. BIVDiff [10] falls short in both content fidelity and temporal consistency, similarly altering background information and character traits during the transfer. AnyV2V [13] excels in maintaining temporal consistency but faces challenges with transfer tasks involving a central object. It often over-transfers the subject’s style, resulting in a loss of its distinctive attributes. The traditional non-diffusion methods, EFDM [38] and CAST [39], compromise detail information during the transfer process and introduce color artifacts. On the contrary, UniVST demonstrates a remarkable ability to balance content fidelity with style richness, achieving a harmonious integration of both. It ensures that the key structural and motion details of the original content are retained, while preserving the primary style of the object with great precision.
We have also compared our method with existing commercial models, namely Gen-1 [51] and Gen-3, as illustrated in Fig.  6. Our approach demonstrates clear superiority in two critical aspects: achieving a closer style similarity to the reference image and maintaining higher content fidelity to the original video. Moreover, it uniquely excels in preserving the style consistency of local objects, ensuring a more cohesive and visually appealing result, even in complex or dynamic scenes.
@@ 5.4 Ablation Studies
We have developed five distinct variants to assess the contribution of its individual components: (1) The full UniVST model, (2) UniVST without mask guidance, (3) UniVST without latent-shift, (4) UniVST without attention-shift, and (5) UniVST without window smoothing. The quantitative outcomes are presented in Table II, while the qualitative results are displayed in Fig. 8.
Our mask-guided strategy effectively maintains the foreground style throughout the transfer process, showing its efficacy both quantitatively and qualitatively, without necessitating additional model components. The latent-shift and attention-shift strategies are both integral to the style transfer process. It is noteworthy noteworthy that the absence of the attention-shift strategy results in an incomplete transfer and a blurring of local details, as exemplified in Fig. 8. Finally, our optical flow smoothing strategy significantly enhances the modeling of temporal consistency, effectively mitigating video artifacts. In addition, we conducted an in-depth study of the three proposed modules on the DAVIS-Laion dataset to further substantiate their individual contributions and overall impact.
Figure 7: Visual results of mask propagation on the DAVIS dataset [42]. The mask is blended with the original image to emphasize the subject, highlighting it while dimming the background.
Figure 8: Visual results for ablations. Various components in our framework have all played important roles.
Figure 9: Quantitative results of mask propagation with different sampling ratio $r$ on the DAVIS [42] dataset. Best view with zooming in.
TABLE III: Performance and efficiency of mask propagation with different anchor frames. The bold fonts represent our final configuration.
@@
Previous-1(Naive) 0.5321 0.6505 2.365
Previous-1+First 0.7799 0.8714 2.423($\downarrow$2.5%)
Previous-3+First 0.7792 0.8711 2.477($\downarrow$4.7%)
Previous-5+First 0.7800 0.8716 2.577($\downarrow$9.0%)
Previous-7+First 0.7809 0.8721 2.673($\downarrow$13.0%)
Previous-9+First 0.7816 0.8726 2.742($\downarrow$15.9%)
Previous-11+First 0.7820 0.8728 2.868($\downarrow$21.3%)
Previous-13+First 0.7820 0.8728 2.873($\downarrow$21.5%)
Previous-15+First 0.7820 0.8729 2.878($\downarrow$21.7%)
TABLE IV: Performance and efficiency of mask propagation with different values of $k$. The bold fonts represent our final configuration.
@@
Mask Propagation. We first present the visualization of the final mask propagation results, as shown in Fig. 7. Then, we further discuss the impact of downsampling rates on propagation accuracy and inference speed using the DAVIS dataset [42]. Detailed results can be found in Fig. 9. As the sampling rate decreases, the efficiency of our mask propagation is significantly improved, while effectively maintaining the accuracy of the final propagation results. We also examine the impact of anchor frames selection on propagation accuracy and efficiency. As shown in Table III, introducing more anchor frames improves propagation accuracy but also results in longer propagation times. Furthermore, we find that as the number of frames increases, the accuracy gains become less significant. After weighing these trade-offs, we select the first frame and the previous 9 frames as the final hyperparameter configuration. In addition, we perform an ablation study on the hyperparameter $k$ of $k$-NN [33]. As shown in Table IV, the highest propagation accuracy is achieved at $k=15$, without affecting propagation efficiency.
TABLE V: Quantitative results of attention-shift with different timestep intervals on the DAVIS-Laion dataset. The bold fonts represent our final configuration.
Timesteps $\leq 0.6T$ $\leq 0.5T$ $\leq 0.4T$ $\leq 0.3T$ $\leq 0.2T$ $\geq\mathbf{0.4T}$ $\geq 0.5T$ $\geq 0.6T$ $\geq 0.7T$ $\geq 0.8T$
@@
TABLE VI: Quantitative results of latent-shift with different timestep intervals on the DAVIS-Laion dataset. The bold fonts show our final configuration.
@@
[…]
TABLE VII: Comparison between attention-shift and some existing training-free stylization methods on the DAVIS-Laion dataset. The bold fonts show our final configuration.
Configuration The key-value replacement The key-value AdaIN Ours with Increasing Ours with Decreasing
@@
Figure 10: Visual results of latent-shift with different timestep intervals on the DAVIS-Laion.
@@
Figure 11: Visual results of attention-shift with different timestep intervals on the DAVIS-Laion. Best view with zooming in.
AdaIN-Guided Style Transfer. We investigate the optimal timestep interval for this method in Table V. It means that applying attention-shift in the early and middle stages of denoising can yield better results. Futhermore, we explore the optimal timestep intervals for latent-shift, with quantitative results provided in Table VI. Applying latent-shift in the late stages yields better results. In addition, we present visual results of these ablation experiments, as shown in Fig. 10 and Fig. 11. For latent-shift, applying it in the early stages of denoising can cause issues such as content blurriness and semantic loss during the transfer process. In contrast, attention-shift produces better visual results when applied in the early stages of denoising than in the later stages. Finally, to further investigate the effectiveness of our proposed attention-shift and the monotonicity of the time coefficient $\beta_{t}$, we compare the performance of four configurations: key-value replacement, key-value AdaIN, ours with increasing time coefficients, and ours with decreasing time coefficients. As shown in Table VII, using a decreasing time coefficient enables attention-shift to outperform other training-free methods.
Figure 12: Qualitative comparisons of different inversion methods on transfer quality. EasyInv [34] demonstrates better transfer quality compared to DDIM-Inv [16].
Figure 13: Visual results of sliding-window consistent smoothing with different window sizes. Best view with zooming in.
Sliding-Window Consistency Smoothing. To assess the impact of sliding window size on editing results and computational costs, we present the visual effects and corresponding inference times for different window sizes in Fig. 13. While larger window sizes yield slight visual improvements, they also result in significant inference costs. On the other hand, we further explore how applying this strategy at different timestep intervals affects final temporal consistency. As shown in Fig. 14, we find that the quality of smoothing strategy varies across the early, middle, and late stages of the denoising process, with the best results achieved when it is applied during the middle stage.
@@
Figure 14: Quantitative results of sliding-window consistent smoothing with varying timestep intervals on the DAVIS-Laion.
Impact of Inversion on UniVST. As a training-free stylization method, we need to obtain the initial noise of the original video, meaning that our method can not bypass inversion technique. Therefore, the performance of this method largely depends on the quality of the initial noise obtained through inversion. In Fig. 12, we find that naive inversion can lead to a loss of semantic details during the transfer process and adversely affect subsequent mask propagation. To minimize the errors introduced by the inversion process, we adopt EasyInv [34] and show better visual effects.
Figure 15: Two examples of multi-step style transfer are demonstrated: the first involves stylizing the background, followed by an additional stylization of the foreground, building upon the initial transformation.
6 Multi-Step Video Style Transfer

### v4

@@ 5.2 Quantitative Comparison
We evaluate the stylized videos generated by different methods across four aspects, in order to comprehensively assess these methods’ performance. (1) Overall transfer performance. Drawing inspiration from the methodologies outlined in [17, 30], we assess the overall transfer performance using ArtFID [48], a comprehensive metric that combines LPIPS [49] for gauging content fidelity and FID [50] for assessing style richness. (2) Local style preservation. To evaluate the preservation of local style, we utilize SSIM [51] metric for the texture preservation, while CLIP-I score [52] is utilized to assess the preservation of semantic content. During the evaluation, the background is neutralized to ensure accuracy. (3) Temporal consistency. In line with the approaches detailed in [30, 11], we measure the similarity between consecutive frames using the CLIP-F score [52], which serves as a measure of inter-frame coherence in the edited video. (4) Inference efficiency. To assess the inference efficiency, we calculate the aggregate time during inference, thereby assessing their performance in terms of computational speed. Furthermore, we employ six established methods as baselines and benchmark them against our UniVST on two datasets: DAVTG-WikiArt and DAVTG-Laion. Quantitative comparisons are comprehensively shown in Table I and Table II.
@@
The results demonstrate that our UniVST outperforms all competing methods across three performance metrics. On both the DAVTG-WikiArt and DAVTG-Laion datasets, UniVST consistently surpasses existing methods in local style preservation and temporal consistency, showcasing its superiority in localized video style transfer tasks over existing methods. Regarding overall transfer performance, UniVST achieves the best results on the DAVTG-Laion dataset and secures the second spot on DAVTG-WikiArt. Localized transfer tasks might lead to a marginally higher FID score, slightly affecting overall performance compared to other methods. Nevertheless, these findings underscore UniVST’s proficiency in effectively balancing content fidelity and style richness.
Our inference efficiency does not stand out significantly. This is because we opt not to adopt the temporal attention mechanisms in existing T2V frameworks [11, 30]. Instead, we utilize the sliding-window consistent smoothing scheme, which requires dynamic optical flow estimation during the subsequent denoising process. While this approach improves temporal consistency, it also leads to increased computational overhead. The detailed discussion on the inference efficiency of each module will be provided in Sec. 5.5.
TABLE III: Ablation study on the DAVTG-WikiArt dataset. Bold values indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Temporal Consistency
@@
W/o all components (baseline) 45.676 31.380 0.411 0.979 0.976 0.981
W/o localized latent blending 38.014 26.050 0.405 0.966 0.970 0.979
W/o AdaIN-guided latent-shift 37.649 26.909 0.349 0.985 0.990 0.980
W/o AdaIN-guided attention-shift 39.589 27.506 0.388 0.985 0.988 0.983
W/o sliding-window consistent smoothing 38.626 26.907 0.384 0.982 0.989 0.970
W/ all components (ours) 37.152 26.134 0.369 0.986 0.991 0.980
@@ 5.3 Qualitative Comparison
For qualitative evaluation, we have juxtaposed the editing results with those of six baseline methods, as illustrated in Fig. 6. Diffutoon [12], which is renowned for its expertise in cartoon coloring, exhibits limited responsiveness to a variety of image styles. This results in a final edited video that is notably monotonous, and heavily inclined towards a cartoonish anime aesthetic. StyleCrafter [11], tailored for style-based video generation, demonstrates a high sensitivity to stylistic cues. However, it encounters difficulties in preserving content fidelity, leading to substantial distortion of background details and character features during the transfer process. BIVDiff [10] falls short in both content fidelity and temporal consistency, similarly altering background information and character traits during the transfer. AnyV2V [13] excels in maintaining temporal consistency but faces challenges with transfer tasks involving a central object. It often over-transfers the subject’s style, resulting in a loss of its distinctive attributes. As for the traditional non-diffusion methods, EFDM [20] and CAST [19], they degrade fine details and induce color artifacts during the style transfer process. On the contrary, our UniVST exhibits a remarkable ability to balance content fidelity and style richness, achieving a harmonious integration of both. It ensures that the key structural and motion details of the original videos are retained, while preserving the primary style of the object with great precision.
TABLE IV: Ablation study on the DAVTG-Laion dataset. Bold values indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Temporal Consistency
[…]
W/o all components (baseline) 38.981 26.594 0.413 0.979 0.976 0.981
W/o localized latent blending 30.848 20.974 0.404 0.961 0.967 0.979
W/o AdaIN-guided latent-shift 31.001 21.766 0.362 0.985 0.989 0.980
W/o AdaIN-guided attention-shift 32.534 22.623 0.377 0.985 0.989 0.981
W/o sliding-window consistent smoothing 30.834 21.249 0.386 0.981 0.988 0.971
W/ all components (ours) 30.636 21.146 0.383 0.986 0.990 0.981
Figure 7: Visual results for ablations. Various components in our framework have all played important roles. Best view with zooming-in.
@@ 5.4 Ablation Studies
In this section, we first validate the effectiveness of each core component in our proposed UniVST. Then, we conduct an in-depth analysis of hyper-parameter settings for the three modules: point-matching mask propagation, AdaIN-guided localized video stylization, and sliding-window consistency smoothing, individually.
TABLE V: Accuracy and efficiency of mask propagation under different value of $n$ on DAVIS-Laion dataset, Bold fonts represent final configuration.
@@
n=1 0.5321 0.6505 2.365 (Base)
n=1 + First 0.7799 0.8714 2.423 ($\downarrow$2.5%)
n=3 + First 0.7792 0.8711 2.477 ($\downarrow$4.7%)
n=5 + First 0.7800 0.8716 2.577 ($\downarrow$9.0%)
n=7 + First 0.7809 0.8721 2.673 ($\downarrow$13.0%)
n=9 + First 0.7816 0.8726 2.742 ($\downarrow$15.9%)
n=11 + First 0.7820 0.8728 2.868 ($\downarrow$21.3%)
n=13 + First 0.7820 0.8728 2.873 ($\downarrow$21.5%)
n=15 + First 0.7820 0.8729 2.878 ($\downarrow$21.7%)
Figure 8: Quantitative results of mask propagation with different sampling ratio $r$ on DAVIS [44] dataset. Best view with zooming-in.
TABLE VI: Accuracy and efficiency of mask propagation with different values of $k$ on DAVIS-Laion dataset. Bold fonts represent final configuration.
@@
5.4.1 Ablation Study on Core Components
We have developed six distinct variants to assess the contribution of its individual components: (1) UniVST without all components (baseline), (2) UniVST without localized latent blending, (3) UniVST without AdaIN-guided latent-shift, (4) UniVST without AdaIN-guided attention-shift, (5) UniVST without sliding-window consistent smoothing and (6) UniVST with all components (ours). The quantitative outcomes are presented in Table III and Table IV, while the qualitative results are displayed in Fig. 7. Our mask-guided strategy effectively preserves the foreground style during the stylization process, demonstrating its efficacy quantitatively and qualitatively, without the need for additional models like SAM [35]. The AdaIN-guided latent-shift and attention-shift strategies are essential for style transfer. Notably, omitting the attention-shift strategy leads to incomplete transfer and blurred local details, as shown in Fig. 7. Finally, our sliding-window consistency smoothing scheme enhances temporal consistency and effectively reduces video artifacts. In the following section, we conduct an in-depth study of the three proposed modules on the DAVIS-Laion dataset to further substantiate their individual contributions, respectively.
Figure 9: Visual results of our point-matching mask propagation strategy on the DAVIS dataset [44]. The masks are blended with the original video to emphasize the foreground, while dimming the background. Our propagation strategy can effectively adapt to various complex scenarios.
TABLE VII: Quantitative results of AdaIN-guided attention-shift with different timestep intervals on DAVIS-Laion dataset. Bold fonts indicate final configuration.
Timestep Intervals $\leq 0.6T$ $\leq 0.5T$ $\leq 0.4T$ $\leq 0.3T$ $\leq 0.2T$ $\geq\mathbf{0.4T}$ $\geq 0.5T$ $\geq 0.6T$ $\geq 0.7T$ $\geq 0.8T$
@@ 5.4.1 Ablation Study on Core Components
TABLE VIII: Quantitative results of AdaIN-guided latent-shift with different timestep intervals on DAVIS-Laion dataset. Bold fonts indicate final configuration.
Timestep Intervals $\mathbf{0.1T}\sim\mathbf{0.2T}$ $0.2T\sim 0.3T$ $0.4T\sim 0.5T$ $0.5T\sim 0.6T$ $0.7T\sim 0.8T$ $0.8T\sim 0.9T$
@@
TABLE IX: Comparison of our method with other training-free stylization methods on DAVIS-Laion dataset. Bold values indicate final configuration.
Method K-V Replacement K-V AdaIN Ours ($\beta_{t}\uparrow$) Ours ($\beta_{t}\downarrow$)
@@
5.4.2 Analysis of Point-Matching Mask Propagation
We first investigate the effect of down-sampling rates on propagation accuracy and inference time using the DAVIS dataset [44]. Detailed results are presented in Fig. 8. As the down-sampling rate $r$ decreases, the efficiency of our mask propagation improves significantly, while maintaining the accuracy of the final propagation results. Next, we also examine the impact of anchor frame selection on propagation accuracy and efficiency. As shown in Table V, increasing the number of anchor frames improves propagation accuracy but also leads to longer propagation times. Moreover, we observe that as the number of frames increases, the accuracy gains diminish. After considering these trade-offs, we choose the first frame and the previous $n=9$ frames as the final hyper-parameter configuration. We also compare the results for different values of the hyper-parameter $k$ in $k$-NN [37]. As shown in Table VI, the highest propagation accuracy is consistently achieved when $k=15$, without sacrificing the overall propagation efficiency.
We present qualitative examples of our mask propagation results in Fig. 9. In the short-term propagation setting, we evaluate scenarios involving fast-moving foreground objects, deformations, and occlusions. Despite these complexities, our method demonstrates precise object tracking and robust mask alignment, consistently preserving accuracy across diverse conditions. For long-term propagation, our approach continues to deliver strong performance. The integration of anchor frames enables stable and accurate mask propagation over extended temporal distances, mitigating typical degradation issues associated with long sequences. However, certain limitations persist, which will be examined in Sec. 5.5.
@@ 5.4.2 Analysis of Point-Matching Mask Propagation
Figure 10: Visual results of latent-shift with different timestep intervals on DAVIS-Laion dataset. Best view with zooming-in.
@@
Figure 11: Quantitative results of sliding-window consistent smoothing with varying timestep intervals. Best view with zooming-in.
5.4.3 Analysis of AdaIN-Guided Localized Video Stylization
We investigate the optimal time interval for this method in Table VII. It means that applying AdaIN-guided attention-shift in the early and middle stages of denoising can yield better results. Furthermore, we explore the optimal timestep intervals for AdaIN-guided latent-shift, with quantitative results provided in Table VIII. Applying AdaIN-guided latent-shift in the late stages yields better results. In addition, we present visual results of these ablation experiments, as shown in Fig. 10 and Fig. 12. For AdaIN-guided latent-shift, applying it in the early stages of denoising can cause issues such as content blurriness and semantic loss during the transfer process. In contrast, AdaIN-guided attention-shift produces better visual results when applied in the early stages of denoising than in the later stages.
Figure 12: Visual results of attention-shift with different timestep intervals on DAVIS-Laion dataset. Best view with zooming-in.
Figure 13: Visual results of sliding-window consistent smoothing with different window sizes on DAVIS-Laion dataset. Best view with zooming-in.
Finally, to further investigate the effectiveness of our proposed AdaIN-guided attention-shift mechanism and the impact of the monotonicity of the time coefficient $\beta_{t}$, we compare the performance of four configurations: K-V replacement [17], K-V AdaIN, our method with increasing time coefficients, and our method with decreasing time coefficients. As shown in Table IX, using decreasing time coefficient enables the AdaIN-guided attention-shift to outperform other training-free methods, such as K-V replacement, K-V AdaIN, or attention-shift with increasing time coefficients.
Figure 14: Inference time distribution of different components in UniVST. Sliding-window consistent smoothing exhibits an inference bottleneck.
5.4.4 Analysis of Sliding-Window Consistency Smoothing
To assess the impact of different window sizes $m$ on stylistic effects and computational costs in the sliding-window consistency smoothing scheme, we present the visual effects and corresponding inference time for various window sizes in Fig. 13. While larger window sizes yield slight visual improvements, they also result in significant inference costs. Ultimately, we select $m=2$ to balance stylistic effects and computational costs. Furthermore, we also explore how applying this strategy at different timestep intervals influences the final temporal consistency. As shown in Fig. 11, we observe that the effectiveness of the smoothing strategy varies across the early, middle, and late stages of the denoising process, with the best results achieved during the middle stage.
Figure 15: Qualitative comparisons of different inversion methods. EasyInv [38] demonstrates better transfer quality compared to DDIM-Inv [16].
5.5 Addtional Analysis
In this section, we further analyze the bottleneck of inference time, the constraints in mask propagation, and the impacts of inversion.
Bottleneck of Inference Time. In Table I and Table II, our proposed UniVST exhibits the longest inference time among all compared methods, revealing a key computational bottleneck. To better understand this limitation, we perform a detailed breakdown of the inference time distribution across the core modules of UniVST, as illustrated in Fig. 14. These modules include: DDIM Inversion, Point-Matching Mask Propagation, AdaIN-Guided Localized Video Stylization, and Sliding-Window Consistent Smoothing. Our analysis indicates that the primary source of inference overhead arises from the sliding-window consistent smoothing module. This component relies on dynamic optical flow estimation during the denoising phase, which improves temporal consistency but also adds significant computational cost. In contrast, most existing text-to-video (T2V) frameworks adopt temporal attention mechanisms, which are generally more efficient though potentially less robust in handling complex motion. To address this trade-off, a promising alternative is to replace the smoothing module with a 3D VAE decoder that directly reconstructs temporally coherent video from the denoised latent representation, thereby maintaining temporal consistency without incurring additional inference time.
Constraints in Mask Propagation. Although our point-matching mask propagation strategy enables zero-shot object localization, as demonstrated in Fig. 9, it presents several limitations: 1. Challenges in Multi-Object Scenarios. In scenes containing multiple similar objects (e.g., two individuals), the semantic similarity between instances can lead to confusion during propagation. This often results in incorrect mask associations, particularly as the propagation distance increases, causing error accumulation over time. While our current evaluation focuses on diverse yet single-object scenes—where such issues have minimal impact—extending the method to robustly handle multi-object scenarios remains an important direction for future work. 2. Limitations in Fine-Grained Mask Propagation. Although the propagation achieves an average Dice score of approximately $0.87$, it occasionally fails to capture fine object boundaries, leading to incomplete or imprecise tracking. As shown in Fig. 9, while the method effectively handles fast motion and occlusions, it struggles to retain detailed edge information. Nonetheless, our localized latent blending strategy, which performs dynamic region-wise fusion during the denoising phase, effectively mitigates this limitation by reducing reliance on high-precision masks. The mechanism can still preserve the object’s stylization quality even without guidance from high-precision masks.
Impacts of Inversion. As a training-free stylization method, UniVST relies on the initial noise derived from the original video, thereby rendering the inversion step crucial. The overall performance of UniVST is somewhat sensitive to the quality of the inverted noise, which serves as a critical foundation for subsequent stylization. Inadequate inversion can compromise both content fidelity and stylistic consistency. As illustrated in Fig. 15, we observe that naive inversion methods often fail to preserve fine-grained semantic details and adversely impact subsequent mask propagation. To mitigate the artifacts introduced by imperfect inversion, we employ EasyInv [38], a more advanced inversion technique. By leveraging EasyInv, we achieve higher-fidelity noise reconstructions, which in turn lead to improved visual quality and stable mask propagation throughout the stylization process.
Figure 17: Multi-step video style transfer: first, the background is stylized, and then the foreground is stylized based on the transformed background.
5.6 Addtional Comparison
Futhermore, we have also compared our UniVST with existing commercial models, namely Gen-1 [53] and Gen-3, as illustrated in Fig. 16. UniVST demonstrates clear superiority in two critical aspects: achieving a closer style similarity to the reference image and maintaining higher content fidelity to the original video. In addition, UniVST uniquely excels in accurately preserving the style consistency of local objects, thereby ensuring a more cohesive and visually appealing result, even in complex or dynamic scenes.
5.7 Multi-Step Video Style Transfer